The Corner Grocer has a 7-year, 6.5 percent semiannual coupon bond outstanding with a $1,000 par value. The bond has a yield to maturity of 5.5 percent. How much will the bond price decrease if the market yield suddenly increases to 7 percent?

Please double check my numbers..

The first bond Second bond

n 14 n 14

I/y 2.75    I/y 3.5

pv ? 1057.50 pv ? 972.70

pmt -32.50    pmt -32.50

fv -1000    fv -1000

Answer is: 84.76 ???
Answer #1

Calculating Bond Price at YTM = 5.5%

Using TVM Calculation,

PV = [FV = 1,000, PMT = 32.50, N = 14, I = 0.055/2]

PV = $1,057.46

Calculating Bond Price at YTM = 7.0%

Using TVM Calculation,

PV = [FV = 1,000, PMT = 32.50, N = 14, I = 0.07/2]

PV = $972.70

Change in Bond Price = 972.70 - 1,057.46

Change in Price = -$84.76